# Finding radio transients with anomaly detection and active learning based on volunteer classifications

### v1

@@ Abstract
In this work we explore the applicability of unsupervised machine learning algorithms to the task of finding radio transients. Facilities such as the Square Kilometre Array (SKA) will provide huge volumes of data in which to detect rare transients; the challenge for astronomers is how to find them. We demonstrate the effectiveness of anomaly detection algorithms using 1.3 GHz light curves from the SKA precursor MeerKAT. We make use of three sets of descriptive parameters (‘feature sets’) as applied to two anomaly detection techniques in the Astronomaly package and analyse our performance by comparison with citizen science labels on the same dataset. By using transients found by volunteers as our ground truth, we demonstrate that anomaly detection techniques can recall over half of the radio transients in the 10 per cent of the data with the highest anomaly scores. We find that the choice of anomaly detection algorithm makes a minimal (sub-one per cent) level difference, but that feature set choice is crucial, especially when considering available resources for human inspection and/or follow-up. Active learning, where human labels are given for just 2 per cent of the data, improves recall by up to 10 percentage points, depending on the feature-model pair. The best performing feature-model pairs result in a factor of 5 times fewer sources requiring vetting by experts. This is the first effort to apply anomaly detection techniques to finding radio transients and shows great promise for application to other datasets, and as a real-time transient detection system for upcoming large surveys.
@@
pubyear: 2024pagerange: Finding radio transients with anomaly detection and active learning based on volunteer classificationsFinding radio transients with anomaly detection and active learning based on volunteer classifications
@@ 1 Introduction
The Square Kilometre Array (SKA; Braun et al., 2019) promises to deliver unprecedented data across many areas of astrophysics. Already, SKA precursors such as ASKAP (Hotan et al., 2021) and MeerKAT (Jonas & MeerKAT Team, 2016) are providing a unique combination of highly sensitive observations taken at regular cadence over a wide field-of-view. These observations then allow for untargeted searches for radio transients in surveys such as ThunderKAT (Fender et al., 2016) and VAST (Variables and Slow Transients, Murphy et al., 2013). ThunderKAT observed known radio transients (e.g. X-ray binaries; XRBs), but the unique capabilities of MeerKAT have allowed for serendipitous discovery of magnetically active and flaring stars (Driessen et al., 2020, 2022a; Andersson et al., 2022; Chastain et al., 2023; Peters, 2023; Fijma et al., 2024, as well as Mlangeni et al. and Nyamai et al., in prep), scintillating AGN and unique pulsar behaviour (Driessen et al., 2022b; Rowlinson et al., 2022). The VAST survey’s wide-field imaging with ASKAP has discovered large, scintillation-inducing plasma arcs (Wang et al., 2021a), mysterious transients near the Galactic centre (Wang et al., 2021b, 2022), and many highly variable systems on minute timescales (Wang et al., 2023). Furthermore, recent observations at timescales of a few minutes have led to uniquely long-period transients, pulsars and white dwarf systems (Hurley-Walker et al., 2022, 2023, 2024; Caleb et al., 2022, 2024; Pelisoli et al., 2023; de Ruiter et al., 2024). Finally, imaging on very short timescales can immediately localise fast radio bursts (FRBs) to high precision (Rhodes et al., 2023; Driessen et al., 2024). The combination of all aspects of a dataset - its spectral coverage, time baseline and polarimetry - has been shown to provide a unique discovery space for pulsars and transients (Heywood, 2023; Smirnov et al., 2024).
@@
Despite the advances in appropriate feature use, anomaly detection and active learning in optical surveys, these tools have yet to be applied to the rapidly-growing field of wide-field radio transient surveys, with the most similar use-cases being for investigating anomalous spectrograms and parameter optimisation in low-frequency radio data and pipelines (Rowlinson et al., 2019; Mesarcik et al., 2023). The Astronomaly package (Lochner & Bassett, 2021) provides a general framework for discovering anomalies in astronomical datasets, which we can use to test these different aspects of the discovery process. Astronomaly has three main scientific steps consisting of feature extraction, anomaly detection and active learning, whilst also providing data management, pre-processing and visualisation tools. Exploration of how these different stages perform, with different features and algorithms applied to a novel dataset form sections 34.3 of this work, though we note that they are all built into Astronomaly’s framework. Astronomaly has been used to discover variable and flaring stars in optical surveys (Webb et al., 2020, 2021), as well as rare and unique morphologies of optical galaxies, mergers and radio sources (Etsebeth et al., 2023; Mohale & Lochner, 2023; Lochner et al., 2023).
@@
We make use of 8874 light curves from the MeerKAT telescope as part of commensal ThunderKAT observations, analysed by citizen scientists in search of variables and transients (Andersson et al., 2023). By its nature, unsupervised learning requires no classifications and therefore easily scales to large datasets as there is no need for costly and/or time-consuming labelling of data nor the training of large models. By the same token, however, it is hard to verify performance of unsupervised learning methods without some known ‘ground truth’ (see e.g. Giles & Walkowicz, 2019). Therefore, we make use of the citizen science classifications as ground truth against which to compare anomaly detection methods from the Astronomaly package. We describe the MeerKAT and citizen science data in section 2 before detailing the feature sets used in this work in section 3. We then demonstrate how standard anomaly detection performs on these data in section 4 and how the active learning built into Astronomaly improves this performance in section 4.3. We discuss this work and its implications before concluding in sections 6 and 7.
@@ 2.1 Observational Data
Once imaged, light curves are extracted from the observations using the Transient Pipeline (TraP, Swinbank et al., 2015). The TraP extracts flux density measurements for all sources detected with a S/N $\geq$8 within 1.5$\times$ the primary beam radius of each observation. Each flux measurement comes from a Gaussian fit to each source, fixed to be the size of the PSF in that observation, as we are only interested in point sources and a changing PSF between epochs can induce false variability. We also note that the TraP is run such that if, in any observation, a source is found, that position is checked and force-fit in all subsequent epochs (but not in prior images), providing upper-limits in images where the source is absent (e.g. a transient). Examples of a bona fide transient, a false positive and a stable light curve are all visible in Figure 1, demonstrating the diversity of our data.
@@ 3 Feature Extraction
In order to make use of our anomaly detection tools, we must first calculate some features that describe our light curves. The choice of how to describe a dataset is crucial in machine learning tasks and can generally be split between features based on domain knowledge or experience and deep learning, where features are learned automatically from the data. In this work we use two domain-informed feature sets which have been used in machine learning searches for variability in previous studies. We choose to use these domain-informed features as opposed to those learnt from the data for several reasons. Firstly, automatic feature extraction methods can have many thousands of parameters (internal weights and biases) to update, requiring similarly extensive datasets. With our moderate sample of $\sim$9,000 light curves, taken from heterogeneously observed fields, we cannot expect our data to be representative of all possible underlying variability. Similarly, data-driven feature extraction can be computationally expensive, requiring many CPU- or GPU-hours to train. Finally, by using features tested in other regimes we can compare and contrast their performance between use cases. We will also compare the performance of the two feature sets to a baseline pair of parameters often used in transient detection and variable characterisation (see e.g. the ThunderKAT commensal work in section 1).
@@ 3.1 Baseline features
As a baseline feature set we make use of two parameters often used to discover radio transients. For a light curve of $N$ flux density measurements $F_{i}\pm\sigma_{i}$, observed at frequency $\nu$, the two variability statistics are defined as
@@
where $s$, $\overline{F_{\nu}}$ and $\overline{F_{\nu}}^{*}$ denote standard deviation, mean and weighted average of the flux density respectively. Broadly speaking, $\eta_{\nu}$, as a chi-squared test, quantifies the statistical significance of variability in a light curve, whilst $V_{\nu}$ describes the amplitude of variability and is sometimes known as the modulation index. We note that this two paramter setup is far simpler (lower dimensions) than the other two feature sets used, which may not capture more subtle variations in the light curves. This is due to our anomaly detection techniques being designed for high dimensional data and there is no guarantee that their algorithmic definitions of anomalies match human intuition when using $\eta_{\nu}$ and $V_{\nu}$.
@@ 3.2 Feets features
The next set of features used comes from the feets package (Cabral et al., 2018a, b), inspired by the FATS code (Nun et al., 2015). This package combines over 40 features used for time series analysis, which can be broken into approximately three subsets. First, some features calculated are simple statistical quantities, such the mean, amplitude and standard deviation of a light curve. The second subset of features were designed for the classification of variable stars (Richards et al., 2011) and quasars (Kim et al., 2011) and include, for example, the skew, kurtoisis and the exponent of a light curve’s structure function. Finally, several features used are derived from a Lomb-Scargle periodogram (Lomb, 1976; Scargle, 1982) and its resultant best fit to each light curve. We note that there are some features in this codebase that require colour information but as our data is univariate (monochromatic) we cannot employ these features in this implementation. These feets features have an implementation built into Astronomaly’s framework, making them ready to test for a given application out-of-the-box. The full table of features used is detailed in Table 3.2.
@@
Skew $\frac{N}{(N-1)(N-2)}\sum(F_{i}-\bar{F}/\sigma)^{3}$ $F_{i}$ Richards et al. (2011)
@@ 3.3 Wavelet features
The second set of features used come from studies primarily aimed at describing and classifying supernovae (Lochner et al., 2016; Alves et al., 2022), though they have also been used in the classification of radio transients (Sooknunan et al., 2021). These features, called wavelet features, arise from the understanding that time series $f(t)$ can be decomposed into a sum of basis functions
@@
with coefficients $a_{i}$, where $\phi_{i}(t)$ might for example be a set of sines and cosines (i.e. a Fourier decomposition). We can then use the coefficients of our basis functions as a set of features for later use. Here, we use the symlet family of wavelets as our basis functions (see e.g. Arfaoui et al., 2021 for more on these), due to their successful use in previous work (Lochner et al., 2016; Alves et al., 2022). These wavelet features are calculated with minimal assumptions on the data and so it is our hope that they describe the observed light curves well, as opposed to the feets features, whose calculations are all made with some idea of canonical variability statistics (e.g. from quasars) in mind. Furthermore, wavelet transforms are translation equivariant and so the same object (e.g. a SN) observed at different times since peak brightness should produced similar coefficients. We calculate the wavelet coefficients for our light curves using the snmachine codebase developed by the LSST Dark Energy Science Collaboration (Lochner et al., 2016; Alves et al., 2022). In order to do this, we first interpolate our light curves with a Gaussian Process Regression (Rasmussen & Williams, 2006), allowing us to account for gaps in our uneven time series. To do this, all light curves are shifted to take place between 00 and $t_{max}$ days, where for our dataset $t_{max}$ = 1300 days, corresponding to our longest light curves in the GX339--4 field. We then evaluate our regression every $\sim$5 days (i.e. at 276 instances), approximately matching our typical observational cadence. Example Gaussian process regression fits can be seen in Figure 1.
@@
Once all light curves are interpolated by the regression, the wavelet transform is calculated, producing a set of 275 highly redundant coefficients. In order to reduce this high dimensionality, we run a Principal Component Analysis (PCA; Pearson, 1901), which transforms the data into an orthogonal basis. In this now-orthogonal data space, we choose to preserve the 40 largest eigenvalues for each light curve, which captures 99.999995 per cent of the summative variance of the data. This results in reducing the data space from 276 features per light curve to only 40, while retaining almost all of the variance. The choice of 40 principal components is a balance between reducing the dimensionality of the features whilst trying to preserving enough information such that the anomalies, which constitute a small minority of all of the light curves, are retained. In a truly unsupervised setting, it is not possible to known how many eigenvalues to preserve as one has no labels to train from.
@@
Finally, it is worth noting that extracting this set of features is the most computationally expensive part of our analysis. However, even the slowest step, the Gaussian process regression, runs in 10 minutes for the entire database on a standard desktop using a 3.5 GHz quad-core processor. Also, all of these features, once calculated, can be saved and these routines then need not be run again. This is an important consideration when considering scalability for larger datasets or for a real-time transient detection system.

### v2

@@ Abstract
In this work we explore the applicability of unsupervised machine learning algorithms to finding radio transients. Facilities such as the Square Kilometre Array (SKA) will provide huge volumes of data in which to detect rare transients; the challenge for astronomers is how to find them. We demonstrate the effectiveness of anomaly detection algorithms using 1.3 GHz light curves from the SKA precursor MeerKAT. We make use of three sets of descriptive parameters (‘feature sets’) as applied to two anomaly detection techniques in the Astronomaly package and analyse our performance by comparison with citizen science labels on the same dataset. Using transients found by volunteers as our ground truth, we demonstrate that anomaly detection techniques can recall over half of the radio transients in the 10 per cent of the data with the highest anomaly scores. We find that the choice of anomaly detection algorithm makes a minor difference, but that feature set choice is crucial, especially when considering available resources for human inspection and/or follow-up. Active learning, where human labels are given for just 2 per cent of the data, improves recall by up to 20 percentage points, depending on the combination of features and model used. The best performing results produce a factor of 5 times fewer sources requiring vetting by experts. This is the first effort to apply anomaly detection techniques to finding radio transients and shows great promise for application to other datasets, and as a real-time transient detection system for upcoming large surveys.
@@
pubyear: 2025pagerange: Finding radio transients with anomaly detection and active learning based on volunteer classificationsFinding radio transients with anomaly detection and active learning based on volunteer classifications
@@ 1 Introduction
The Square Kilometre Array (SKA; Braun et al., 2019) promises to deliver unprecedented data across many areas of astrophysics. Already, SKA precursors such as ASKAP (Hotan et al., 2021) and MeerKAT (Jonas & MeerKAT Team, 2016) are providing a unique combination of highly sensitive observations taken at regular cadence over a wide field-of-view. These observations then allow for untargeted searches for radio transients in surveys such as ThunderKAT (Fender et al., 2016) and VAST (Variables and Slow Transients, Murphy et al., 2013). ThunderKAT observed known radio transients (e.g. X-ray binaries; XRBs), but the unique capabilities of MeerKAT have allowed for serendipitous discovery of magnetically active and flaring stars (Driessen et al., 2020, 2022a; Andersson et al., 2022; Chastain et al., 2023, 2024; Peters, 2023; Fijma et al., 2024; Smirnov et al., 2025, as well as Mlangeni et al. and Nyamai et al., in prep), scintillating AGN and unique pulsar behaviour (Driessen et al., 2022b; Rowlinson et al., 2022). The VAST survey’s wide-field imaging with ASKAP has discovered large, scintillation-inducing plasma arcs (Wang et al., 2021a), mysterious transients near the Galactic centre (Wang et al., 2021b, 2022), and many highly variable systems on minute timescales (Wang et al., 2023). Furthermore, recent observations at timescales of a few minutes have led to uniquely long-period transients [^1] Here, minute-timescales are considered ‘long’ with respect to typical pulsar periods of $<1$ s., pulsars and white dwarf systems (Hurley-Walker et al., 2022, 2023, 2024; Caleb et al., 2022, 2024; Pelisoli et al., 2023; de Ruiter et al., 2024). Finally, imaging on very short timescales can immediately localise fast radio bursts (FRBs) to high precision (Rhodes et al., 2023; Driessen et al., 2024). The combination of all aspects of a dataset - its spectral coverage, time baseline and polarimetry - has been shown to provide a unique discovery space for pulsars and transients (Heywood, 2023; Smirnov et al., 2024).
@@
Despite the advances in appropriate feature use, anomaly detection and active learning in optical surveys, these tools have yet to be applied to the rapidly-growing field of wide-field radio transient surveys, with the most similar use-cases being for investigating anomalous spectrograms and parameter optimisation in low-frequency radio data and pipelines (Rowlinson et al., 2019; Mesarcik et al., 2023). The Astronomaly package (Lochner & Bassett, 2021) provides a general framework for discovering anomalies in astronomical datasets, which we can use to test these different aspects of the discovery process. Astronomaly has three main scientific steps consisting of feature extraction, anomaly detection and active learning, whilst also providing data management, pre-processing and visualisation tools. Exploration of how these different stages perform, with different features and algorithms applied to a novel dataset form sections 34.3 of this work, though we note that they are all built into Astronomaly’s framework. Astronomaly has been used to discover variable and flaring stars in optical surveys (Webb et al., 2020, 2021), as well as rare and unique morphologies of optical galaxies, mergers and radio sources (Etsebeth et al., 2024; Mohale & Lochner, 2024; Lochner et al., 2023).
@@
We make use of 8874 light curves from the MeerKAT telescope as part of commensal ThunderKAT observations, analysed by citizen scientists in search of variables and transients (Andersson et al., 2023). By its nature, unsupervised learning requires no classifications and therefore easily scales to large datasets as there is no need for costly and/or time-consuming labelling of data nor the training of large models. By the same token, however, it is hard to verify performance of unsupervised learning methods without some known ‘ground truth’ (see e.g. Giles & Walkowicz, 2019). Therefore, we make use of the citizen science classifications as ground truth against which to compare anomaly detection methods from the Astronomaly package. We describe the MeerKAT and citizen science data in section 2 before detailing the features used in this work in section 3. We then demonstrate how standard anomaly detection performs on these data in section 4 and how the active learning built into Astronomaly improves this performance in section 4.3. We discuss this work and its implications before concluding in sections 6 and 7.
@@ 2.1 Observational Data
Once imaged, light curves are extracted from the observations using the Transient Pipeline (TraP, Swinbank et al., 2015). The TraP extracts flux density measurements for all sources detected with a S/N $\geq$8 within 1.5$\times$ the primary beam radius of each observation. Each flux measurement comes from a Gaussian fit to each source, fixed to be the size of the PSF in that observation, as we are only interested in point sources and a changing PSF between epochs can induce false variability. The TraP is run such that if, in any observation, a source is found, that position is checked and force-fit in all subsequent epochs (but not in prior images), providing upper-limits in images where the source is absent (e.g. a transient). Examples of a bona fide transient, a false positive and a stable light curve are all visible in Figure 4, demonstrating the diversity of our data.
@@ 3 Feature Extraction
In order to make use of our anomaly detection tools, we must first calculate some features that describe our light curves. The choice of how to describe a dataset is crucial in machine learning tasks and can generally be split between features based on domain knowledge or experience and deep learning, where features are learned automatically from the data. In this work we use two domain-informed feature sets which have been used in machine learning searches for variability in previous studies. We also test the union of both these sets together, which we will refer to as the combined feature set. We choose to use these domain-informed features as opposed to those learnt from the data for several reasons. Firstly, automatic feature extraction methods can have many thousands of parameters (internal weights and biases) to update, requiring similarly extensive datasets. With our moderate sample of $\sim$9,000 light curves, taken from heterogeneously observed fields, we cannot expect our data to be representative of all possible underlying variability. Similarly, data-driven feature extraction can be computationally expensive, requiring many CPU- or GPU-hours to train. Finally, by using features tested in other regimes we can compare and contrast their performance between use cases. We will also compare the performance of the two feature sets to a baseline pair of parameters often used in transient detection and variable characterisation (see e.g. the ThunderKAT commensal work in section 1).
@@ 3.1 Baseline features
As a baseline pair of features we make use of two parameters often used to discover radio transients. For a light curve of $N$ flux density measurements $F_{i}\pm\sigma_{i}$, observed at frequency $\nu$, the two variability statistics are defined as
@@
where $s$, $\overline{F_{\nu}}$ and $\overline{F_{\nu}}^{*}$ denote standard deviation, mean and weighted average of the flux density respectively. Broadly speaking, $\eta_{\nu}$, as a chi-squared test, quantifies the statistical significance of variability in a light curve, whilst $V_{\nu}$ describes the amplitude of variability and is sometimes known as the modulation index. We note that this two parameter setup is far simpler (lower dimensions) than the two feature sets described subsequently, which may not capture more subtle variations in the light curves. This is due to our anomaly detection techniques being designed for high dimensional data and there is no guarantee that their algorithmic definitions of anomalies match human intuition when using $\eta_{\nu}$ and $V_{\nu}$.
@@ 3.2 Feets features
Next we used a feature set from the feets package (Cabral et al., 2018a, b), inspired by the FATS code (Nun et al., 2015). This package combines over 40 features used for time series analysis, which can be broken into approximately three subsets. First, some features calculated are simple statistical quantities, such as the mean, amplitude, and standard deviation of a light curve. The second subset of features used are included due to their known statistical or astrophysical relevance to a given class of object such as variable stars (Richards et al., 2011) and quasars (Kim et al., 2011) and include, for example, the skewness and the exponent of a light curve’s structure function. Finally, several features used are derived from a Lomb-Scargle periodogram (Lomb, 1976; Scargle, 1982) and its resultant best fit to each light curve. We note that there are some features in this codebase that require colour information but as our data is univariate (monochromatic) we cannot employ these features in this implementation. These feets features have an implementation built into Astronomaly’s framework, making them ready to test for a given application out-of-the-box. The full table of features used is detailed in Table 3.2.
@@
Skewness $\frac{N}{(N-1)(N-2)}\sum(F_{i}-\bar{F}/\sigma)^{3}$ $F_{i}$ Richards et al. (2011)
@@ 3.3 Wavelet features
The final set of features used come from studies primarily aimed at describing and classifying supernovae (Lochner et al., 2016; Alves et al., 2022), though they have also been used in the classification of radio transients (Sooknunan et al., 2021). These features, called wavelet features, arise from the understanding that time series $f(t)$ can be decomposed into a sum of basis functions
@@
with coefficients $a_{i}$, where $\phi_{i}(t)$ might for example be a set of sines and cosines (i.e. a Fourier decomposition). We can then use the coefficients of our basis functions as a set of features for later use. Here, we use the symlet family of wavelets as our basis functions (see e.g. Arfaoui et al., 2021 for more on these), due to their successful use in previous work (Lochner et al., 2016; Alves et al., 2022). These wavelet features are calculated with minimal assumptions on the data and so it is our hope that they describe the observed light curves well, whilst by contrast the feets package’s parametric calculations are made with assumptions on the statistical distributions of the data. Furthermore, wavelet transforms are translation equivariant and so the same object (e.g. a SN) observed at different times since peak brightness should produced similar coefficients.
We calculate the wavelet coefficients for our light curves using the snmachine codebase developed by the LSST Dark Energy Science Collaboration (Lochner et al., 2016; Alves et al., 2022). In order to do this, we first interpolate our light curves with a Gaussian Process Regression (GP; Rasmussen & Williams, 2006), allowing us to account for gaps in our uneven time series. To do this, all light curves are shifted to take place between 00 and $t_{max}$ days, where for our dataset $t_{max}$ = 1300 days, corresponding to our longest light curves in the GX339--4 field. The underlying procedure built into snmachine fits Gaussian processes from the George library [^6] https://george.readthedocs.io/en/latest/, with this implementation using GPs with a mean of zero and one of two covariance kernels. The two kernels tested are an exponential square kernel of the form
$k_{\mathrm{exp}}(t_{i},t_{j})=\mathrm{exp}\left(-\frac{|t_{i}-t_{j}|^{2}}{2}\right),$ (5)
and a combination of $k_{\mathrm{exp}}(t_{i},t_{j})$ with an exponential sine squared kernel of scale parameter $\Gamma$ and period $P$ i.e.
$k_{\mathrm{sin}}(t_{i},t_{j})=k_{\mathrm{exp}}(t_{i},t_{j})+\mathrm{exp}\left(-\Gamma\sin^{2}\left[\frac{\pi}{P}|t_{i}-t_{j}|\right]\right).$ (6)
These two kernels are intended to capture any periodic behaviour whilst the exponential decays account for the idea that points separated further in time are expected to be less correlated. The kernel used is determined by minimising the negative log-likelihood of the GP with the SciPy implementation of a limited-memory Broyden–Fletcher–Goldfarb–Shanno algorithm (Byrd et al., 1995; Virtanen et al., 2020). Once the Gaussian Process has been fit, we evaluate the regression every $\sim$5 days (i.e. at 276 instances), approximately matching our typical observational cadence. Example Gaussian process regression fits can be seen in Figure 4. For light curves shorter than $t_{max}$ in length, the remaining datapoints are assumed to be 0.
@@
Once all light curves are interpolated by the regression, the wavelet transform is calculated, producing a set of 275 highly redundant coefficients. In order to reduce this high dimensionality, we run a Principal Component Analysis (PCA; Pearson, 1901), which transforms the data into an orthogonal basis. In this now-orthogonal data space, we choose to preserve the 40 largest eigenvalues for each light curve, which captures 99.999995 per cent of the summative variance of the data. This results in reducing the data space from 276 features per light curve to only 40, while retaining almost all of the variance. Fewer principal components would result in faster downstream computations, however as mentioned earlier, anomalies (transients) constitute a small minority of our sample whilst the majority are stable light curves which should require few wavelet features to describe. Furthermore, light curves shorter than $t_{max}$ are assumed to be 0 past their final datapoint, which will contribute towards the majority of the dataset being non-variable. As a result, initial testing with only 10 principal components showed markedly worse performance. This shows that the diversity of light curve lengths affects the resultant wavelets and their principal components retained, which may bias results. The choice of 40 principal components is a balance between reducing the dimensionality of the features whilst trying to preserving enough information such that the anomalies, which constitute a small minority of all of the light curves, are retained. In a truly unsupervised setting, it is not possible to know how many eigenvalues to preserve as one has no labels to train from.
@@
Finally, it is worth noting that extracting this set of features is the most computationally expensive part of our analysis. However, even the slowest step, the Gaussian process regression, runs in 10 minutes for the entire database on a standard desktop using a 3.5 GHz quad-core processor. Also, all of these features, once calculated, can be saved and these routines then need not be run again, making static comparisons like this work fast to conduct. For a real-time system features may be re-calculated for every new datapoint, however this may not be optimal or required. For example, the Gaussian Process Regressions that underlie the wavelet features scale naïvely as $\mathcal{O}(N^{3})$ and so calculating them for each new datapoint may be prohibitive. For larger datasets or for a real-time transient detection system a balance between computational scalability and information gained may be required and could dictate what feature sets can be deployed for a given science case.